# Spin flip locking by the tunneling and relaxation in a driven double quantum dot with spin-orbit coupling

### v1

@@ Abstract
Coupled spin evolution and tunneling together with the relaxation and decoherence effects are studied for the double quantum dot formed in a semiconductor nanowire and driven by the periodic electric field. Such system represents a model of the spin and charge qubits interacting via the strong spin-orbit coupling. It is found that at certain regimes the combination of fast relaxation in the coordinate channel with the slower relaxation in the spin channel leads to the promising combination of fast spin manipulation and slow spin relaxation, locking the flipped spin in an excited state in one of the dots for a sufficiently long time. The predicted effect is maintained for a wide range of the relaxation times and the driving amplitude both for the coordinate and the spin channels and is also observed on higher subharmonic which requires lower driving frequencies.
@@ I Introduction
The spin dynamics in low dimensional semiconductor structures demonstrates a great variety of both fundamental phenomena and promising applications in spintronics and quantum computations. The basic building block representing the qubit, i.e. the two-level system, has been the subject of intensive research where different phenomena such as the Landau-Zener-Stückelberg-Majorana (LZSM) interference Nori2010 ; Nori2023 have been studied. The coupled tunneling and the spin evolution in semiconductor structures with quantum dots have been the subject of intensive research Semenov2007 ; Taylor2007 ; Platero2008 ; Petersson2010 ; Ludwig2013 ; Ribeiro2013 ; Vandersypen2013 ; Vandersypen2014 ; Loss2019 ; Krzywda2021 ; Platero2023 including the experimentally observed manifestation of the interplay between the tunneling and the spin flip associated with the electric dipole spin resonance (EDSR) and its subharmonics in the gate-defined single and double quantum dots (QD) in semiconductor structures Loss2002 ; Rashba2003 ; Loss2003 ; Debald2005 ; Loss2006 ; Yokoshi2009 ; Nowack2007 ; StehlikPetta2014 ; Studenikin2018 ; Studenikin2019 ; Studenikin2023 . Such systems may represent a model of the hybrid spin-charge qubit Petta2006 ; Burkard2007 ; Zhang2007 ; Ramon2010 ; Petersson2010b ; Mizuta2017 ; Yang2019 ; Kregar2020 ; Rancic2022 where the Zeeman splitting generates the two-level subspace of the spin subsystem and the double or multi-dot Platero2022 confinement potential provides a controlled charge distribution for the two-level charge (coordinate) subspace encoding the states by the electron location in left- or right-QD. The two qubits may interact via the spin-orbit coupling intrinsically present in the given structure or generated by the non-uniform magnetic field PioroLadriere2007 and driven by the monochromatic or two-photon Raman Bodey2019 mechanism.
@@
Motivated by the mentioned experiments, we have studied in our recent papers KS2022 ; PRB2023 the LZSM effects for the coupled spin and charge dynamics driven in the double quantum dots by the periodic electric field and the spin-orbit coupling (SOC), focusing on the EDSR KS2022 and its subharmonics PRB2023 . It was found that the tunneling may enhance the SOC-induced spin flip via the interdot tunneling. The effects studied in these papers has been considered on relatively short time scales which allowed to neglect the relaxation effects and to consider only the coherent dynamics.
@@
When several, at least two interacting quantum dots (or qubits) are considered on the time interval comparable with the relaxation time in any of the subsystems, the enlarged size of the structure and the associated transport and tunneling effects inevitably call for the consideration of various relaxation and decoherence effects studied under the well-known density matrix formalism BP ; Hanggi1997 ; GrifoniHanggi1998 ; Hone2009 ; Nalbach2009 ; Xu2014 . The applications include the already mentioned low dimensional semiconductor structures with quantum dots Semenov2007 ; Taylor2007 ; Platero2008 ; Petersson2010 ; Ludwig2013 ; Vandersypen2013 ; Vandersypen2013 ; Krzywda2021 ; Platero2023 , the qubits based on the Josephson junctions Makhlin2001 ; Satanin2012 ; Bastrakova2021 ; Pashin2023 , the circuit or cavity quantum electrodynamics Bonifacio2020 ; Tokman2023 , the entangled photons Marques2015 , the three-level systems Zhang2023 , and many others. The primary question is how well the results obtained within the coherent dynamics approximation survive on the specific time intervals and what qualitatively new effects the relaxation and decoherence may introduce into the coupled tunneling and spin dynamics in addition to the trivial damping and smearing of the valuable signals.
@@
In this paper we include the relaxation and decoherence within the master equation approach for the electron density matrix and study the driven spin dynamics coupled with the interdot tunneling in a double quantum dot under the periodic driving by the electric field. Similar to our previous studies KS2022 ; PRB2023 , we consider the resonant spin-flip transitions assisted by the interdot tunneling. It is found that under some circumstances the relaxation can play an unexpectedly stabilizing role for the spin flip, locking the flipped spin in the excited Zeeman state. The mechanism is activated in the resonant regime where the spin flip is accompanied with the interdot tunneling. The fast charge relaxation drives the electron into the single dot having the lower energy for the charge degree of freedom while the spin maintains its flipped state on the longer time scale which may go beyond the spin relaxation time due to the driving. This effect, similar to the ones observed in three-level atomic and laser systems Evers2005 ; Ilinova2012 ; Boukhris2023 but observed mainly in the $GHz$ frequency range in the semiconductor dot experiments Nowack2007 ; StehlikPetta2014 ; Studenikin2018 ; Studenikin2019 ; Studenikin2023 , can be observed for various system parameters and both on the main and the sub-harmonics of the resonance, is of fundamental interest and is promising for the spin manipulation technique development in nanostructures. Our predictions may stimulate the development of coupled spin and charge qubit setups where the charge dynamics improves the characteristics of the spin evolution.
@@
This paper is organized as follows. In Sec. II we describe the Hamiltonian and its lowest four-level subspace forming the coupled spin and charge qubits. The degree of the validity for the four-level model for the coherent dynamics is evaluated by searching for the borders in the parameter space which enclose the impact of the four lowest levels above the preset threshold. We extend the approach by introducing the evolution equation for the electron density matrix with the Lindblad operators and the relaxation rates for the charge and spin two-level subsystems, respectively. In Sec. III we present the numerical analysis of the regimes where the spin flip is assisted by the interdot tunneling and visualize the spin evolution on the associated Bloch sphere for each qubit subsystem. We consider different types of resonances both without and with the interdot tunneling. In the regime with the interdot tunneling, we observe the unexpectedly stabilizing role of the relaxation at certain regimes on the spin flip in a single dot where the electron has been initialized in the ground state. Finally, in Sec. IV we give our conclusions.
@@ II.1 The Hamiltonian
We may call the Zeeman-split pair of states (,)($\downarrow$,$\uparrow$)( ↓ , ↑ ) in each dot as the spin subsystem (or spin qubit) with the distance between the levels $\Delta_{1}$ shown on the left panel of Fig.1.
@@
where the numerical calculation of the energy levels gives us $\Delta_{2}=\gamma|U_{d}|$ with $\gamma\approx 0.55$. For the negative detuning $U_{d}<0$ with $|U_{d}|\ll U_{0}$ the ground state $\varepsilon_{1}$ with the wavefunction $\phi_{1}(x)$ is located in the right QD with the exponentially small overlap with the left QD and the next state $\phi_{2}(x)$ is located in the left QD. This makes the basis for the charge two-level subsystem by attributing the wavefunction location to the right and left QD, respectively. If we restrict ourselves to this two-level subspace for the coordinate degree of freedom we may call it as representation of a charge qubit which is shown schematically on the right panel in Fig.1.
The spin and charge qubits are coupled via the third term in (2) which is the SOC that may contain both Dresselhaus and Rashba terms. Here we will limit ourselves to the Dresselhaus term only since the Rashba term provides similar impact on the spin and charge evolution PRB2023 . For the leading order in the wavevector the SOC is linear for the GaAs-based nanowire,
@@
and obtain the set of energy levels $E_{n}$ and the two-component eigenfunctions $\Psi_{n}(x)$. The principal scheme for the lowest four-level subspace in (8) describing the two coupled spin and charge qubits is given in the center panel of Fig.1 for the negative detuning $U_{d}<0$. In this basic example the level pairs ($E_{1}$, $E_{2}$) as well as ($E_{3}$, $E_{4}$) represent the spin qubit with the splittings in each dot
@@
which are slightly different from each other and from the Zeeman splitting $\Delta_{1}$ from (5) due to the corrections from the SOC term (7). The level pairs ($E_{1}$, $E_{3}$) and ($E_{2}$, $E_{4}$) belonging to the neighboring dots represent the charge qubit with the splitting
@@
where again due to SOC the splittings $\Delta^{(1,2)}_{2}$ are slightly different from the spinless level splitting $\Delta_{2}$ defined in (6). It should be noted that due to the presence of the SOC the amplitude of the $z$-projection of the spin along the direction of the magnetic field for an eigenstate of the Hamiltonian (2) is always less than $1$. It means that these eigenstates in (8) are not pure spin-up or spin-down states, although for the chosen parameters their in-plane spin projection is sufficiently smaller (about $1\%$ in magnitude) than the $z$-projection. This allows us the consideration of the spin-split states in each QD as the states of the single spin subsystem (spin qubit) similar to the pure Zeeman-splitted states (,)($\downarrow$,$\uparrow$)( ↓ , ↑ ).
Figure 1: Scheme of the spin (left) and charge (right) qubits in a double quantum dot with the lowest subspace of the levels $E_{1},\ldots,E_{4}$ of the coupled spin-charge qubit (center). The splitting $\Delta_{1}$ for the spin qubit is generated by the Zeeman term (5), the splitting $\Delta_{2}$ for the charge qubit is generated by the detuning which shifts the relative position of the dot energy minimum. The two qubits are coupled via SOC and driven by the periodic electric field $V_{d}(t)$.
@@ II.2 Four-level approximation
First, we will take a look on the coherent driven dynamics without the relaxation in order to determine the regions in the parameter space where the four-level approximation is valid with the preset tolerance which is needful for the model of two coupled qubits described above. The usual solution for the time-dependent Schrödinger equation $i\partial\psi/\partial t=H\psi$ with the Hamiltonian (1) is found as a sum of the eigenfunctions from (8) with time-dependent coefficients KS2022 ; PRB2023 :
@@
In Fig.2 the black, grey and white dashed lines indicate the resonance transitions between the levels $E_{1}-E_{4}$, $E_{1}-E_{3}$ and $E_{1}-E_{2}$ in Fig.1 for the energy splitting parameters in (5) and (6) taken as $\Delta_{1}=20\mu eV$ and $\Delta_{2}=55\mu eV$, and the dotted lines of the same colors indicate their second subharmonics. From Fig.2 one can observe the leakage probability of $10^{-8}\ldots 10^{-7}$ for all operating frequencies in our modeling which provides a good justification of working within the lowest four-level subspace for the most parts of the parameter map.
A remarkable feature in Fig.2 is that the leakage probability is enhanced at certain frequencies, i.e. on some horizontal lines in Fig.2 where and obvious increase is seen also with the growing driving amplitude. These frequencies, grouped at $\omega\sim 20\mu eV$ being the spin qubit splitting (9), correspond to the the hybridization of the $E_{1}-E_{2}$ level resonance and the high ($k\sim 100$) subharmonic of the $E_{1}-E_{5}$ resonance with the interlevel distance $E_{5}-E_{1}\sim 2000\mu eV$. Such amplification of the resonance at the hybridization points in the parameter space have been observed in our preceding paper PRB2023 . We can draw a conclusion that operating on the basic EDSR frequency $\omega=E_{2}-E_{1}$ seems to be not the best option due to the intensified leakage to the higher levels above the four-level manifold. A better way off flipping the spin can be the operations on the $E_{1}-E_{4}$ resonance or its subharmonic labeled by the black dashed and dotted lines in Fig.2, respectively. We will gather more justifications on this choice in the next Section.
II.3 Two-qubit representation
As it can be seen from Fig.2, most of the driving field parameter plane provides a good background for the modeling the evolution within the four-level ground subspace $E_{1},\ldots,E_{4}$ illustrated by Fig.1 which is spanned by the wavefunctions $\Psi_{1},\ldots\Psi_{4}$ obtained from (8). We will consider them as the states for the two-qubit representation of our system, the first qubit being the spin qubit on the pair of levels $E_{1}$-$E_{2}$ or $E_{3}$-$E_{4}$ with the level spacing (9) and the second qubit is the charge qubit on the levels $E_{1}$-$E_{3}$ or $E_{2}$-$E_{4}$ with the level spacing (10) provided by the detuning. The two qubits interact via the SOC (7) and are both driven by the periodic field (11).
The description of the Hamiltonian for the two interacting qubits can be made in terms of the Pauli matrices $\sigma_{k}^{(q)}$, $k=0,\ldots,3$, attributed to each of the spin ($q=1$) and charge ($q=2$) qubits:

### v2

@@ Abstract
Coupled spin evolution and tunneling together with the relaxation and decoherence effects are studied for the double quantum dot formed in a semiconductor nanowire and driven by the periodic electric field. Such structure represents a model of the spin and charge subsystems interacting via the strong spin-orbit coupling. It is found that at certain regimes the combination of fast relaxation in the coordinate channel with the slower relaxation in the spin channel leads to the promising combination of fast spin manipulation and slow spin relaxation, locking the flipped spin in an excited state in one of the dots for a sufficiently long time. The predicted effect is maintained for a wide range of the relaxation times and the driving amplitude both for the coordinate and the spin channels and is also observed on higher subharmonic which requires lower driving frequencies.
@@ I Introduction
The spin dynamics in low dimensional semiconductor structures demonstrates a great variety of both fundamental phenomena and promising applications in spintronics and quantum computations. The basic building block representing the qubit, i.e. the two-level system, has been the subject of intensive research where different phenomena such as the Landau-Zener-Stückelberg-Majorana (LZSM) interference Nori2010 ; Nori2023 have been studied. The coupled tunneling and the spin evolution in semiconductor structures with quantum dots have been the subject of intensive research Semenov2007 ; Taylor2007 ; Platero2008 ; Petersson2010 ; Ludwig2013 ; Ribeiro2013 ; Vandersypen2013 ; Vandersypen2014 ; Loss2019 ; Krzywda2021 ; Platero2023 including the experimentally observed manifestation of the interplay between the tunneling and the spin flip associated with the electric dipole spin resonance (EDSR) and its subharmonics in the gate-defined single and double quantum dots (QD) in semiconductor structures Loss2002 ; Rashba2003 ; Loss2003 ; Debald2005 ; Loss2006 ; Yokoshi2009 ; Nowack2007 ; StehlikPetta2014 ; Studenikin2018 ; Studenikin2019 ; Studenikin2023 . Such systems may represent a model of the hybrid spin-charge qubit encoded in the the spin and charge interacting subsystems Petta2006 ; Burkard2007 ; Zhang2007 ; Ramon2010 ; Petersson2010b ; Mizuta2017 ; Yang2019 ; Kregar2020 ; Rancic2022 . Here the Zeeman splitting generates the two-level subspace of the spin subsystem and the double or multi-dot Platero2022 confinement potential provides a controlled charge distribution for the two-level charge (coordinate) subspace encoding the states by the electron location in left- or right-QD. The two subsystems may interact via the spin-orbit coupling intrinsically present in the given structure or generated by the non-uniform magnetic field PioroLadriere2007 and driven by the monochromatic or two-photon Raman Bodey2019 mechanism. In a recent paper Zuk2024 the spin locking mechanism has been proposed as a tool for the dynamic decoupling in a system of the transmon qubits aimed on the reduction of the gate errors. Thus, it is of interest to consider further examples of the interacting spin and charge subsystems where the spin flip locking may take place prolonging the lifetime of the excited state.
@@
Motivated by the mentioned experiments, we have studied in our recent papers KS2022 ; PRB2023 the LZSM effects for the coupled spin and charge dynamics driven in the double quantum dots by the periodic electric field and the spin-orbit coupling (SOC), focusing on the EDSR KS2022 and its subharmonics PRB2023 . It was found that the tunneling may enhance the SOC-induced spin flip via the interdot tunneling. The effects studied in these papers have been considered on relatively short time scales which allowed to neglect the relaxation effects and to consider only the coherent dynamics.
@@
When several, at least two interacting quantum dots are considered on the time interval comparable with the relaxation time in any of the subsystems, the enlarged size of the structure and the associated transport and tunneling effects inevitably call for the consideration of various relaxation and decoherence effects studied under the well-known density matrix formalism BP ; Hanggi1997 ; GrifoniHanggi1998 ; Hone2009 ; Nalbach2009 ; Xu2014 . The applications include the already mentioned low dimensional semiconductor structures with quantum dots Semenov2007 ; Taylor2007 ; Platero2008 ; Petersson2010 ; Ludwig2013 ; Vandersypen2013 ; Vandersypen2013 ; Krzywda2021 ; Platero2023 , the qubits based on the Josephson junctions Makhlin2001 ; Satanin2012 ; Bastrakova2021 ; Pashin2023 , the circuit or cavity quantum electrodynamics Bonifacio2020 ; Tokman2023 , the entangled photons Marques2015 , the three-level systems Zhang2023 , and many others. The primary question is how well the results obtained within the coherent dynamics approximation survive on the specific time intervals and what qualitatively new effects the relaxation and decoherence may introduce into the coupled tunneling and spin dynamics in addition to the trivial damping and smearing of the valuable signals.
@@
In this paper we include the relaxation and decoherence within the master equation approach for the electron density matrix and study the driven spin dynamics coupled with the interdot tunneling in a double quantum dot under the periodic driving by the electric field. Similarly to our previous studies KS2022 ; PRB2023 , we consider the resonant spin-flip transitions assisted by the interdot tunneling. It is found that, under some circumstances, the relaxation can play an unexpectedly stabilizing role for the spin flip, locking the flipped spin in the excited Zeeman state. The mechanism is activated in the resonant regime where the spin flip is accompanied by the interdot tunneling. The fast charge relaxation drives the electron into the single dot having the lower energy for the charge degree of freedom while the spin maintains its flipped state on the longer time scale which may go beyond the spin relaxation time due to the driving. This effect, similar to the ones observed in three-level atomic and laser systems Evers2005 ; Ilinova2012 ; Boukhris2023 but observed mainly in the $GHz$ frequency range in the semiconductor dot experiments Nowack2007 ; StehlikPetta2014 ; Studenikin2018 ; Studenikin2019 ; Studenikin2023 , can manifest itself for various system parameters and both on the main and the sub-harmonics of the resonance. It is of fundamental interest and is promising for the spin manipulation technique development in nanostructures. Our predictions may stimulate the development of coupled spin and charge qubit setups where the charge dynamics improves the characteristics of the spin evolution.
@@
This paper is organized as follows. In Sec. II we describe the Hamiltonian and its lowest four-level subspace forming the coupled spin and charge subsystems. The degree of the validity for the four-level model for the coherent dynamics is evaluated by searching for the borders in the parameter space which enclose the impact of the four lowest levels above the preset threshold. We extend the approach by introducing the evolution equation for the electron density matrix with the Lindblad operators and the relaxation rates for the charge and spin two-level subsystems, respectively. In Sec. III we present the numerical analysis of the regimes where the spin flip is assisted by the interdot tunneling and visualize the spin evolution on the associated Bloch sphere for each subsystem. We consider different types of resonances both without and with the interdot tunneling. In the regime with the interdot tunneling, we observe the unexpectedly stabilizing role of the relaxation at certain regimes on the spin flip in a single dot where the electron has been initialized in the ground state. Finally, in Sec. IV we give our conclusions.
@@ II.1 The Hamiltonian
We may call the Zeeman-split pair of states (,)($\downarrow$,$\uparrow$)( ↓ , ↑ ) in each dot as the spin subsystem with the distance between the levels $\Delta_{1}$ shown on the left panel of Fig.1.
@@
where the numerical calculation of the energy levels gives us $\Delta_{2}=\gamma|U_{d}|$ with $\gamma\approx 0.55$. For the negative detuning $U_{d}<0$ with $|U_{d}|\ll U_{0}$ the ground state $\varepsilon_{1}$ with the wavefunction $\phi_{1}(x)$ is located in the right QD with the exponentially small overlap with the left QD and the next state $\phi_{2}(x)$ is located in the left QD. This makes the basis for the charge two-level subsystem by attributing the wavefunction location to the right and left QD, respectively. In the following we will restrict ourselves to this two-level subspace for the coordinate degree of freedom shown schematically on the right panel in Fig.1.
The spin and charge subsystems are coupled via the third term in (2) which is the SOC that may contain both Dresselhaus and Rashba terms. Here we will limit ourselves to the Dresselhaus term only since the Rashba term provides similar impact on the spin and charge evolution PRB2023 . For the leading order in the wavevector the SOC is linear for the GaAs-based nanowire,
@@
and obtain the set of energy levels $E_{n}$ and the two-component eigenfunctions $\Psi_{n}(x)$. The principal scheme for the lowest four-level subspace in (8) describing the two coupled spin and charge subsystems is given in the center panel of Fig.1 for the negative detuning $U_{d}<0$. In this basic example the level pairs ($E_{1}$, $E_{2}$) as well as ($E_{3}$, $E_{4}$) represent the spin subsystem with the splittings in each dot
@@
which are slightly different from each other and from the Zeeman splitting $\Delta_{1}$ from (5) due to the corrections from the SOC term (7). The level pairs ($E_{1}$, $E_{3}$) and ($E_{2}$, $E_{4}$) belonging to the neighboring dots represent the charge subsystem with the splitting
@@
where again due to SOC the splittings $\Delta^{(1,2)}_{2}$ are slightly different from the spinless level splitting $\Delta_{2}$ defined in (6). It should be noted that due to the presence of the SOC the amplitude of the $z$-projection of the spin along the direction of the magnetic field for an eigenstate of the Hamiltonian (2) is always less than $1$. It means that these eigenstates in (8) are not pure spin-up or spin-down states, although for the chosen parameters their in-plane spin projection is sufficiently smaller (about $1\%$ in magnitude) than the $z$-projection. This allows us the consideration of the spin-split states in each QD as the states of the single spin subsystem similar to the pure Zeeman-splitted states (,)($\downarrow$,$\uparrow$)( ↓ , ↑ ).
Figure 1: Scheme of the spin (left) and charge (right) subsystems in a double quantum dot with the lowest subspace of the levels $E_{1},\ldots,E_{4}$ of the coupled spin-charge system (center). The splitting $\Delta_{1}$ for the spin subsystem is generated by the Zeeman term (5), the splitting $\Delta_{2}$ for the charge subsystem is generated by the detuning which shifts the relative position of the dot energy minimum. The two subsystems are coupled via SOC and driven by the periodic electric field $V_{d}(t)$.
@@ II.2 Four-level approximation
First, we will take a look on the coherent driven dynamics without the relaxation in order to determine the regions in the parameter space where the four-level approximation is valid with the preset tolerance which is needful for the model of two coupled subsystems described above. The usual solution for the time-dependent Schrödinger equation $i\partial\psi/\partial t=H\psi$ with the Hamiltonian (1) is found as a sum of the eigenfunctions from (8) with time-dependent coefficients KS2022 ; PRB2023 :
@@
In Fig.2 the black, gray and white dashed lines indicate the resonance transitions between the levels $E_{1}-E_{4}$, $E_{1}-E_{3}$ and $E_{1}-E_{2}$ in Fig.1 for the energy splitting parameters in (5) and (6) taken as $\Delta_{1}=20\mu eV$ and $\Delta_{2}=55\mu eV$, and the dotted lines of the same colors indicate their second subharmonics. From Fig.2 one can observe the leakage probability of $10^{-8}\ldots 10^{-7}$ for all operating frequencies in our modeling which provides a good justification of working within the lowest four-level subspace for the most parts of the parameter map.
A remarkable feature in Fig.2 is that the leakage probability is enhanced at certain frequencies, i.e. on some horizontal lines in Fig.2 where and obvious increase is seen also with the growing driving amplitude. These frequencies, grouped at $\omega\sim 20\mu eV$ being the spin subsystem splitting (9), correspond to the hybridization of the $E_{1}-E_{2}$ level resonance and a high ($k\sim 100$) subharmonic of the $E_{1}-E_{5}$ resonance with the interlevel distance $E_{5}-E_{1}\sim 2000\mu eV$. Such amplification of the resonance at the hybridization points in the parameter space has been observed in our preceding paper PRB2023 . We can draw a conclusion that operating on the basic EDSR frequency $\omega=E_{2}-E_{1}$ seems to be not the best option due to the intensified leakage to the higher levels above the four-level manifold. A better way of flipping the spin can be the operations on the $E_{1}-E_{4}$ resonance or its subharmonic labeled by the black dashed and dotted lines in Fig.2, respectively. We will gather more justifications on this choice in the next Section.
II.3 Two-subsystem representation
As it can be seen from Fig.2, most of the driving field parameter plane provides a good background for the modeling the evolution within the four-level ground subspace $E_{1},\ldots,E_{4}$ illustrated by Fig.1 which is spanned by the wavefunctions $\Psi_{1},\ldots\Psi_{4}$ obtained from (8). We will consider them as the states for the two-subsystem representation of our system, the first one being the spin subsystem and the second one is the charge subsystem. The two subsystems interact via the SOC (7) and are both driven by the periodic field (11). We would like to stress that in our double dot system both the spin and charge subsystems are not localized within a single dot, as it is routinely expected. The double dot setup serves as a source of a four-level manifold describing the ensemble of two coupled dots. The spin and charge subsystems can be extracted after the representation of the time-dependent Hamiltonian (1) for the two interacting parts is made in terms of the Pauli matrices $\sigma_{k}^{(q)}$, $k=0,\ldots,3$, attributed to each of the spin ($q=1$) and charge ($q=2$) subsystems: